OpenStudy (loser66):

Construct a function f: R--> R that is differentiable at x =-1,0,1 and at no other point Please, help

OpenStudy (loser66):

If f is diff. at x =-1,0,1, f is continuous at those points Hence the piecewise function will be constructed under the form of \(f(x) = \begin{cases}h(x)~~if~~~x\in A\\g(x)~~if~~x\in ~~~B\end{cases}\)

Nnesha (nnesha):

\(\huge\color{green}{{\rm BEST }~\rm of~luck !! :)}\)

OpenStudy (zarkon):

try looking at the following for starters \[f(x) = \begin{cases}0~~if~~~x\in \mathbb{R}\backslash\mathbb{Q}\\x^2~~if~~x\in\mathbb{Q}\end{cases}\]